OpenStudy (mayankdevnani):

Last question :

OpenStudy (mayankdevnani):

Which of the following polynomials below is identical to the polynomial \[(x^2-1)(x^2-4)\]

OpenStudy (mayankdevnani):

options are : a) \[x^4+4\] b) \[x^2-5x+4\] c) \[(x^2+1)(x^2+4)\] d) \[(x^2+x-2)(x^2-x-2)\]

hartnn (hartnn):

what do u think ? just by looking ? no solving anything..

OpenStudy (ghazi):

one of your options has got no cubic term and it is similar to that of your given equation ..expand your given equation first

OpenStudy (ghazi):

\[x^4-5x^2+4=0\] does any of your option look like this..after the expansion

OpenStudy (anonymous):

WELL........ TWO POLYNOMIALS ARE IDENTICAL IF THE DIFFERENCE BETWEEN THE POLYNOMIALS IS 0

OpenStudy (mayankdevnani):

\[(x^2-1)=(x-1)^2\] is this right?

hartnn (hartnn):

nopes

OpenStudy (mayankdevnani):

what means nopes

OpenStudy (mayankdevnani):

what nopes??

OpenStudy (mayankdevnani):

@igbasallote

OpenStudy (ghazi):

expand every option and you'll have your answer

OpenStudy (mayankdevnani):

but it was right? you did'nt say??

OpenStudy (anonymous):

NO...... (x^2 -1) and (x-1)^2 are not identical

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

since \[(x ^{2}-1)(x ^{2}-4)=((x -1)(x +1)(x-2)(x+2))\] and u know that option d is equal to \[(x+2)(x-1)(x-2)(x+1)\] what does it implies?

OpenStudy (mayankdevnani):

\[x^2-1^2\]

OpenStudy (anonymous):

BECAUSE (x^2-1) -(x-1)^2 =x^2-1-(x^2-2x+1) =x^2-1-x^2+2x-1 =2x-2 THE DIFFERENCE BETWEEN THE POLYNOMIALS IS NOT O

OpenStudy (mayankdevnani):

@REMAINDER

OpenStudy (anonymous):

no ,how did u get that?

OpenStudy (anonymous):

@mayankdevnani did u understand @REMAINDER

OpenStudy (mayankdevnani):

no

OpenStudy (anonymous):

HE FACTORED ALL THE TERMS OF ALL OPTIONS AND THE QUESTION

OpenStudy (anonymous):

wat i mean is that by sorting option d the one i've simplified u'll have (x-1)(x+1)(x-2)(x+2) which is equal to the factor of the \[(x ^{2}-1)(x ^{2}-4)\]

OpenStudy (mayankdevnani):

\[answer \] is \[x^2+2x^2+1\]

OpenStudy (anonymous):

check this  do u understand this part?